# Temporal Cooperative Games

### v3

@@ 3 The SeqShare Class of Mechanisms
In this section, we introduce a class of reward sharing mechanisms that we later show to uniquely satisfy OIR, I4OA, and SE. As a precursor to this class, we first define two conditions called the basis conditions for a TCG $\langle N,v\rangle$.
@@
$\begin{split}\textstyle(i)\quad&\textstyle\sum_{i\in P(\pi)}x_{i}\geqslant v(\pi),\forall\pi\in\Pi,\\ \textstyle(ii)\quad&\textstyle\sum_{i\in P(\pi^{*}(v))}x_{i}=v(\pi^{*}(v)).\end{split}$ (1)
@@
The importance of the basis condition can be understood from the following result that shows that this condition is necessary for two desirable properties of TCGs. The basis conditions are analogous to the core of classical cooperative games. Certain desirable properties are contingent on these sets to be non-empty, while it has a chance of being empty too.
@@
Proof.
Hence $\phi(\pi^{*})$ satisfies basis conditions. ∎
Remark.
The result above does not guarantee existence of a solution concept satisfying I4OA and SE if a basis solution exists. In the following example we show examples of solutions that satisfy the basis and use the basis solution as the reward for $\pi^{*}$ but still fall short of at least one of the given properties. Existence of a basis solution does not guarantee that trivial solution concepts can satisfy the $3$ properties together, as we illustrate in the following example. This shows that method we propose needs to do a non-trivial work to ensure all three.
@@
Consider the TCG as shown in Figure˜1. It is easy to check that the game has a basis solution, but the reward share $\phi$ given below does not satisfy the OIR property.
12v(1)=1v(1)=1v(12)=2v(12)=2
21v(2)=1v(2)=1v(21)=3v(21)=3
Figure 1: A game that satisfies basis conditions.
In this example, $\pi^{*}=21$, $x_{1}=1,x_{2}=2$. Now, consider the following solution concepts.
@@
Our next endeavor is to construct a solution concept that will satisfy both and OIR.
Our desired class of solution concepts start with the basis conditions to check if a solution exists. If the basis conditions hold, Algorithm˜1 yields a class of solution concepts that satisfies OIR, I4OA, and SE. It is a class because there are multiple possible solution concept $\phi$ that can be given by this algorithm. However, we will show that each of them satisfies the three above desirable properties, and any solution concept satisfying the three is a valid solution concept given by this algorithm. The class SeqShare performs the following steps:
@@
    It computes the marginal contribution made by $\pi(2)=j$, and divides it among $i$ and $j$ such that their rewards $\phi_{i}(ij)$ and $\phi_{j}(ij)$ lies within the upper and lower bounds set for them. The lower bounds for the agents $i\text{ and }j$ are then updated to $\phi_{i}(ij)$ and $\phi_{j}(ij)$ respectively for every sequence containing $ij$ as a prefix.
@@
Figure˜2 provides a visual representation of SeqShare. The non-uniqueness of the solution concept given by SeqShare has two distinct sources. First, the basis solution may not be unique as Equation˜1 may be satisfied by multiple $x$ vectors. Second, given a basis solution, there may be multiple functions that are suitable candidates for $\phi$ as the Improvize function returns a non-unique $y$ vector. Hence, SeqShare is a class based on these two freedom of choice. Once a solution concept $\phi$ is obtained via SeqShare, it can be implemented in an online fashion. This is because $\phi$ gives a reward share for every player in every sequence $\pi\in\Pi$. Hence, whenever an agent appears in any sequence, $\phi$ has a reward share for that agent at that point of arrival.
@@
2:Find $\pi^{*}(v)=\operatorname*{\arg\!\max}_{\pi}v(\pi)$
3:Let $x=\textsc{Basis}(N,v,\pi^{*}(v))$
4:if $x$ = null then
5:  Output null, exit
6:else
7:  $\phi(\pi^{*}(v))=x$
8:  $\phi(\pi)=\mathbf{0},\ \forall\pi\neq\pi^{*}(v)$
9:end if
10:for $i\in N$ do
11:  $\phi_{i}(\{i\})=v(\{i\})$
12:end for
13:for $k=2,\ldots,n$ do
14:  for $\pi\in\Pi,\ |\pi|=k-1$ do
15:    for $j\in N\setminus P(\pi)$ do
16:     $y=$
17:    $\textsc{Improvize}(j,\pi,\phi(\pi),\phi(\pi^{*}(v)))$
18:     $\phi_{i}(\pi+j)=\phi_{i}(\pi)+y_{i},\forall i\in P(\pi)$
19:     $\phi_{j}(\pi+j)=y_{j}$
20:    end for
21:  end for
22:end for
23:Output: $\phi$
@@
2:  if solution to Equation˜1 exists then
3:    return $x$ of Equation˜1 w.r.t. $\pi^{*}(v)$
4:  else
5:    return null
6:  end if
7:end procedure
@@
Figure 2: (a) The first player arrives and all the value generated is assigned to player P1. (b) The second player P2 arrives and the marginal value generated by the arrival of this player is distributed among P1 and P2 in a way that ensures the total reward of each agent is less than the upper bound. The total reward for any agent is the sum of the reward in the previous round and the share of marginal increase assigned in the current round. (c) The third player P3 arrives and the marginal value generated by P3 is distributed among P1, P2 and P3 keeping total rewards lower than the upper bound. (d) Finally, P4 arrives and the marginal value generated by this player can only be divided among P1,P2 and P4 as P3 has already reached its maximum possible reward.
Remark.
While we use a basis solution as part of the algorithm, it should be noted that the existence of a basis solution is a property of a given TCG $\langle N,v\rangle$ and not that of a solution concept.
@@ 3.1 Properties of SeqShare
In this section, we present the results of the properties of SeqShare. We first show that the existence of a basis solutions is sufficient to ensure a solution that satisfies OIR, I4OA, and SE using the procedures of SeqShare. Further, we show that if a game has a solution that satisfies OIR, I4OA, and SE it is a solution concept in the SeqShare class. We have already proved Lemma˜1 that states that the existence of a basis solution is necessary to satisfy I4OA and SE. The following theorem completes this lemma with the OIR property.
@@
Given a TCG $\langle N,v\rangle$ a reward function $\phi$ satisfies OIR, I4OA, and SE iff it is from the class SeqShare.
Proof.
First, observe that for all unit length sequences, the allocation given by Algorithm˜1 trivially satisfies SE. Property I4OA is satisfied due to the fact that $\phi_{i}(\{i\})=v(\{i\})\leqslant\phi_{i}(\pi^{*}(v)),\forall i\in N$, given by the Basis conditions. OIR is trivial too, since the game is monotone and $\phi_{i}(\emptyset)=0,\forall i\in N$.
Hence, by the induction hypothesis, assume that the forward implication of the theorem is true for all $\pi$ such that $|\pi|=k-1,k\geqslant 2$. We show that the implication is true for every $\pi+i$ of length $k$ as well. Consider the function Improvize. For every $\pi$ of length k1k-1 and the existing partial share $\phi(\pi)$, Improvize allocates the reward to agents in such a way that it is at least $\phi_{i}(\pi)$ (required by OIR) and at most $\phi_{i}(\pi^{*})$ (required by I4OA). This allocation is always possible since by the monotonicity of the game ensures $v(\pi+i)\geqslant v(\pi)$ and $v(\pi+i)\leqslant v(\pi^{*})$ (by definition of $\pi^{*}$). Hence, I4OA and OIR are satisfied for each $i\in N$. Also, the construction of the $y$ vectors (reward shares) is such that the sum of rewards of all agents $j\in\pi+i$ equals $v(\pi+i)$, satisfying SE.
Now, set, $\forall i\in P(\pi_{\textup{longshort}})$
$=\phi_{i}(\pi_{\textup{longshort}}).$
@@ 4 Properties Inspired by Shapley in TCGs
Classical cooperative games and the classic solution concept due to Shapley operates on the setting where the characteristic function $v$ depends on a coalition. In TCG, the function $v$ depends on a sequence of players. To compare the solution concepts with that of the Shapley properties for a TCG $\langle N,v\rangle$, we first need to distinguish the solution concepts where it depends on the sequence $\pi$ (we will call this solution concepts as we did so far in this paper) and the ones where it does not depend on the sequence (we will call this extended solution concepts in the rest of this paper). The space of solution concepts is thus given by $\Phi=\{\phi:\Pi\times V\rightarrow\mathbb{R}^{n}\}$, and the space of extended solution concepts is given by $\Psi=\{\psi:V\rightarrow\mathbb{R}^{n}\}$, where $V$ is the set of all monotone characteristic functions. A solution concept, however, can be reduced to an extended solution concept by averaging it over all possible full sequences, i.e., $\pi:P(\pi)=N$, as follows.
@@
Given a TCG $\langle N,v\rangle$ and a solution concept $\phi\in\Phi$, the reduction of $\phi$ to an extended solution concept $\overline{\phi}\in\Psi$ is defined as
$\overline{\phi}_{i}=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi).$ (2)
In parity with the Shapley properties, we will now define the additivity, efficiency and null player in the context of both types of solution concepts. To distinguish them, the properties for the solution concept that depends on the sequence are called sequential properties, while those for the extended solution concept are called extended properties. We begin with the sequential properties. The sequential efficiency property is already defined in Definition˜2.
@@
A solution concept $\phi\in\Phi$ is sequentially additive if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$,
$\phi_{i}(\pi,u)+\phi_{i}(\pi,v)=\phi_{i}(\pi,u+v),\ \forall\pi\in\Pi,\forall i\in N.$ (3)
@@
A solution concept $\phi\in\Phi$ satisfies sequential null player if for every TCG $\langle N,v\rangle$
$v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\phi_{i}(\pi)=0,\ \forall\pi\in\Pi.$ (4)
@@
An extended solution concept $\psi\in\Psi$ satisfies extended efficiency if for every TCG $\langle N,v\rangle$
$\sum_{i\in N}\psi_{i}(v)=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi).$ (5)
@@
An extended solution concept $\psi\in\Psi$ satisfies extended null player if for every TCG $\langle N,v\rangle$
$v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\psi_{i}(v)=0.$ (6)
Note that this definition of extended null player is stronger than that in (Sánchez and Bergantiños, 1997), where a player is null only if it does not improve the worth of a sequence if it is inserted in any position of the sequence, i.e., starting from the first, second, till the last position, while our null player only requires the agent to be added at the end. Since the ‘if’ condition of our definition is weaker, the property ENP is stronger than their null player property. Their characterization of the solution concepts require four (and not three) properties, including symmetry, while we need only three, as we will soon see in Section˜5. We show in the following example that even if we use their definitions of the Shapley properties in this setting, the TCG-appropriate properties of solution concepts are in conflict with that of the Shapley-inspired properties, which is the fundamental conclusion of our paper.
@@
Consider the TCG given in Figure˜3. The only solution that lies in SeqShare is $\phi(1)=(1,0),\phi(2)=(0,1),\phi(12)=(1,0),\phi(21)=(1,1)$. The reduction of this solution is $\overline{\phi}=(1,0.5)$. The solution characterized by Sánchez and Bergantiños (1997) for this game is $\psi=(0.75,0.75)$. This tells us that there exist games for which no $\phi\in\textsc{SeqShare}{}$ can be reduced to the solution given by Sánchez and Bergantiños (1997).
12v(1)=1v(1)=1v(12)=1v(12)=1
21v(2)=1v(2)=1v(21)=2v(21)=2
Figure 3: SeqShare is not compatible with Sánchez and Bergantiños (1997)
@@
An extended solution concept $\psi\in\Psi$ satisfies extended additivity if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$
$\psi_{i}(u)+\psi_{i}(v)=\psi_{i}(u+v),\ \forall i\in N.$ (7)
@@
A solution concept $\phi\in\Phi$ satisfies sequential symmetry if for every TCG $\langle N,v\rangle$, the following holds:
$v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\phi_{i}(\pi)=\phi_{j}(\pi_{\textup{swap}(i,j)}),\forall\pi\in\Pi.$ (8)
@@
An extended solution concept $\psi\in\Psi$ satisfies extended symmetry if for every TCG $\langle N,v\rangle$, the following holds:
$v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\psi_{i}=\psi_{j}.$ (9)

### v4

@@ 3 The SeqShare Class of Mechanisms
In this section, we introduce a class of reward-sharing mechanisms that we later show to uniquely satisfy OIR, I4OA, and SE. As a precursor to this class, we first define two conditions called the basis conditions for a TCG $\langle N,v\rangle$.
@@
$\begin{split}\textstyle(i)\sum_{i\in P(\pi)}x_{i}\geqslant v(\pi),\forall\pi\in\Pi,\ \ (ii)\sum_{i\in P(\pi^{*}(v))}x_{i}=v(\pi^{*}(v)).\end{split}$ (1)
@@
The importance of the basis condition can be understood from the following result that shows that this condition is necessary for two desirable properties of TCGs. The basis conditions are analogous to the core of classical cooperative games. Certain desirable properties are contingent on these sets being non-empty, while it has a chance of being empty too.
@@
Proof Sketch.  Given a TCG $\langle N,v\rangle$ and a solution concept $\phi$ satisfying I4OA and SE, choosing $x=\phi(\pi^{*})$ meets the basis conditions. Condition (ii) follows directly from SE applied to $\pi^{*}$. Condition (i) follows from I4OA, which bounds $\phi_{i}(\pi)$ by $\phi_{i}(\pi^{*})$ for all $i$ and $\pi$, and SE ensures $v(\pi)=\sum_{i}\phi_{i}(\pi)$, implying the required inequality. $\square$
Remark. The result above does not guarantee the existence of a solution concept satisfying I4OA and SE if a basis solution exists. In the following example, we show examples of solutions that satisfy the basis and use the basis solution as the reward for $\pi^{*}$, but still fall short of at least one of the given properties. Existence of a basis solution does not guarantee that trivial solution concepts can satisfy the $3$ properties together, as we illustrate in the following example. This shows that the method we propose needs non-trivial effort to ensure all three.
@@
Consider the TCG: $v(1)=1,v(12)=2,v(2)=1,v(21)=3$. It is easy to check that the game has a basis solution. In this example, $\pi^{*}=21$, $x_{1}=1,x_{2}=2$. Now, consider the following solution concepts.
@@
Our next endeavor is to construct a solution concept that will satisfy all of them.
Our desired class of solution concepts starts with the basis conditions to check if a solution exists. If the basis conditions hold, Algorithm˜1 yields a class of solution concepts that satisfy OIR, I4OA, and SE. It is a class because there are multiple possible solution concepts $\phi$ that can be given by this algorithm. However, we will show that each of them satisfies the three above desirable properties, and any solution concept satisfying the three is a valid solution concept given by this algorithm. The class SeqShare performs the following steps:
@@
    It computes the marginal contribution made by $\pi(2)=j$, and divides it among $i$ and $j$ such that their rewards $\phi_{i}(ij)$ and $\phi_{j}(ij)$ lie within the upper and lower bounds set for them. The lower bounds for the agents $i\text{ and }j$ are then updated to $\phi_{i}(ij)$ and $\phi_{j}(ij)$ respectively for every sequence containing $ij$ as a prefix.
@@
Figure˜1 provides a visual representation of SeqShare. The non-uniqueness of the solution concept given by SeqShare has two distinct sources. First, the basis solution may not be unique as Equation˜1 may be satisfied by multiple $x$ vectors. Second, given a basis solution, there may be multiple functions that are suitable candidates for $\phi$ as the Improvize function returns a non-unique $y$ vector. Hence, SeqShare is a class based on these two freedoms of choice. Once a solution concept $\phi$ is obtained via SeqShare, it can be implemented in an online fashion. This is because $\phi$ gives a reward share for every player in every sequence $\pi\in\Pi$. Hence, whenever an agent appears in any sequence, $\phi$ has a reward share for that agent at that point of arrival.
@@
2:Find $\pi^{*}(v)=\operatorname*{\arg\!\max}_{\pi}v(\pi)$, $x=\textsc{Basis}(N,v,\pi^{*}(v))$
3:if $x$ = null then Output null, exit
4:else $\phi(\pi^{*}(v))=x$; $\phi(\pi)=\mathbf{0},\ \forall\pi\neq\pi^{*}(v)$
5:end if
6:for $i\in N$ do $\phi_{i}(\{i\})=v(\{i\})$
7:end for
8:for $k=2,\ldots,n$ do
9:  for $\pi\in\Pi,\ |\pi|=k-1$ do
10:    for $j\in N\setminus P(\pi)$ do
11:     $y=\textsc{Improvize}(j,\pi,\phi(\pi),\phi(\pi^{*}(v)))$
12:     $\phi_{i}(\pi+j)=\phi_{i}(\pi)+y_{i},\forall i\in P(\pi)$; $\phi_{j}(\pi+j)=y_{j}$
13:    end for
14:  end for
15:end for
16:Output: $\phi$
@@
2:  if solution to Equation˜1 exists then return $x$ of Equation˜1 w.r.t. $\pi^{*}(v)$
3:  else return null
4:  end if
5:end procedure
@@
Figure 1: (a) The first player arrives, and all the value generated is assigned to player P1. (b) The second player P2 arrives, and the marginal value generated by the arrival of this player is distributed among P1 and P2 in a way that ensures the total reward of each agent is less than the upper bound. The total reward for any agent is the sum of the reward in the previous round and the share of the marginal increase assigned in the current round. (c) The third player P3 arrives, and the marginal value generated by P3 is distributed among P1, P2, and P,3 keeping total rewards lower than the upper bound. (d) Finally, P4 arrives, and the marginal value generated by this player can only be divided among P1,P2, and P4 as P3 has already reached its maximum possible reward.
Remark. While we use a basis solution as part of the algorithm, it should be noted that the existence of a basis solution is a property of a given TCG $\langle N,v\rangle$ and not that of a solution concept.
@@ 3.1 Properties of SeqShare
In this section, we present the results of the properties of SeqShare. We first show that the existence of a basis solution is sufficient to ensure a solution that satisfies OIR, I4OA, and SE using the procedures of SeqShare. Further, we show that if a game has a solution that satisfies OIR, I4OA, and SE, it is a solution concept in the SeqShare class. Lemma˜1 states that the existence of a basis solution is necessary to satisfy I4OA and SE. The following theorem completes this lemma with the OIR property.
@@
Given a TCG $\langle N,v\rangle$, a reward function $\phi$ satisfies OIR, I4OA, and SE iff it is from the class SeqShare.
Proof Sketch.  If the game $\langle N,v\rangle$ violates the basis conditions, then by Lemma˜1 no solution concept can satisfy I4OA and SE, and SeqShare returns null. Hence, assume the basis conditions hold.
We first prove the reverse direction constructively using Algorithm˜1. The algorithm initializes the allocation using Basis on the maximal sequence $\pi^{*}(v)$ and assigns zero elsewhere. The construction proceeds by induction on sequence length.
For sequences of length one, SE holds trivially, I4OA follows from the basis conditions, and OIR holds by monotonicity. Assume the properties hold for all sequences of length k1k-1. For a sequence $\pi+i$ of length $k$, the function Improvize allocates the marginal reward so that payoffs are non-decreasing (OIR), bounded above by $\phi(\pi^{)}$ (I4OA), and sum to $v(\pi+i)$ (SE). Thus, the properties hold for all sequences.
For the forward direction, suppose there exists a solution concept $\phi$ satisfying I4OA, SE, and OIR but $\phi\notin\textsc{SeqShare}{}$. For all $\phi^{\prime}\in\textsc{SeqShare}{}$, let $\pi^{\prime}$ be the shortest sequence where $\phi^{\prime}$ and $\phi$ differ, among these choose the $\phi^{\prime}$ that corresponds to the longest such sequence. Let $\pi_{\textup{longshort}}$ be a shortest sequence where this $\phi^{\prime}$ and $\phi$ differ.
Using the hypothesis that $\phi$ satisfies OIR, I4OA, and Efficiency, we can construct a solution in SeqShare that matches $\phi$ on all sequences up to and including $|\pi_{\textup{longshort}}|$, contradicting the choice of $\pi_{\textup{longshort}}$. Hence, every solution concept satisfying I4OA, SE, and OIR belongs to SeqShare. $\square$
@@ 4 Properties Inspired by Shapley in TCGs
Classical cooperative games and the classic solution concept due to Shapley operates on the setting where the characteristic function $v$ depends on a coalition. In TCG, the function $v$ depends on a sequence of players. To compare the solution concepts with that of the Shapley properties for a TCG $\langle N,v\rangle$, we first need to distinguish the solution concepts where it depends on the sequence $\pi$ (we will call this solution concepts as we did so far in this paper) and the ones where it does not depend on the sequence (we will call this extended solution concepts in the rest of this paper). The space of solution concepts is thus given by $\Phi=\{\phi:\Pi\times V\rightarrow\mathbb{R}^{n}\}$, and the space of extended solution concepts is given by $\Psi=\{\psi:V\rightarrow\mathbb{R}^{n}\}$, where $V$ is the set of all monotone characteristic functions. The sets $\Phi$ and $\Psi$ are therefore the order-sensitive and order-oblivious classes of solution concepts, respectively. A solution concept, however, can be reduced to an extended solution concept by averaging it over all possible full sequences, i.e., $\pi:P(\pi)=N$, as follows.
@@
Given a TCG $\langle N,v\rangle$ and a solution concept $\phi\in\Phi$, the reduction of $\phi$ to an extended solution concept $\overline{\phi}\in\Psi$ is defined as $\overline{\phi}_{i}=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}\phi_{i}(\pi).$
In parity with the Shapley properties, we will now define the additivity, efficiency, and null player in the context of both types of solution concepts. To distinguish them, the properties for the solution concept that depends on the sequence are called sequential properties, while those for the extended solution concept are called extended properties. We begin with the sequential properties. The sequential efficiency property is already defined in Definition˜2.
@@
A solution concept $\phi\in\Phi$ is sequentially additive if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$, $\phi_{i}(\pi,u)+\phi_{i}(\pi,v)=\phi_{i}(\pi,u+v),\ \forall\pi\in\Pi,\forall i\in N.$
@@
A solution concept $\phi\in\Phi$ satisfies sequential null player if for every TCG $\langle N,v\rangle$ $v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\phi_{i}(\pi)=0,\ \forall\pi\in\Pi.$
@@
An extended solution concept $\psi\in\Psi$ satisfies extended efficiency if for every TCG $\langle N,v\rangle$ $\sum_{i\in N}\psi_{i}(v)=\frac{1}{n!}\sum_{\pi\in\Pi:P(\pi)=N}v(\pi).$
@@
An extended solution concept $\psi\in\Psi$ satisfies extended null player if for every TCG $\langle N,v\rangle$ $v(\pi+i)=v(\pi),\ \forall\pi\in\Pi_{-i}\implies\psi_{i}(v)=0.$
Note that this definition of extended null player is stronger than that in (Sánchez and Bergantiños, 1997), where a player is null only if it does not improve the worth of a sequence if it is inserted in any position of the sequence, i.e., starting from the first, second, till the last position, while our null player only requires the agent to be added at the end. Since the ‘if’ condition of our definition is weaker, the property ENP is stronger than its null player property. Their characterization of the solution concepts requires four (and not three) properties, including symmetry, while we need only three, as we will soon see in Section˜5. We show in ˜2 that even if we use their definitions of the Shapley properties in this setting, the TCG-appropriate properties of solution concepts are in conflict with those of the Shapley-inspired properties, which is the fundamental conclusion of our paper.
@@
Consider the TCG given by $v(1)=1,v(12)=1,(2)=1,v(21)=2$. The only solution that lies in SeqShare is $\phi(1)=(1,0),\phi(2)=(0,1),\phi(12)=(1,0),\phi(21)=(1,1)$. The reduction of this solution is $\overline{\phi}=(1,0.5)$. The solution characterized by Sánchez and Bergantiños (1997) for this game is $\psi=(0.75,0.75)$. This tells us that there exist games for which no $\phi\in\textsc{SeqShare}{}$ can be reduced to the solution given by Sánchez and Bergantiños (1997).
@@
An extended solution concept $\psi\in\Psi$ satisfies extended additivity if for every pair of TCGs $\langle N,v\rangle$ and $\langle N,u\rangle$ $\psi_{i}(u)+\psi_{i}(v)=\psi_{i}(u+v),\ \forall i\in N.$
@@
A solution concept $\phi\in\Phi$ satisfies sequential symmetry if for every TCG $\langle N,v\rangle$, the following holds $\forall\pi\in\Pi$ $v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\phi_{i}(\pi)=\phi_{j}(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi.$
@@
An extended solution concept $\psi\in\Psi$ satisfies extended symmetry if for every TCG $\langle N,v\rangle$, $v(\pi)=v(\pi_{\textup{swap}(i,j)}),\ \forall\pi\in\Pi\implies\psi_{i}=\psi_{j}.$